OpenStudy (anonymous):

How does (2+ e^(2t) +(-e)^(-2t) )^.5 = e^t+e^-t work?

OpenStudy (anonymous):

how does it simplify

OpenStudy (jamesj):

Check and see what\[(e^t + e^{-t})^2\] equals and you'll see what is going on here.

OpenStudy (jamesj):

(and the equation you've written down isn't quite right, as you'll see)

OpenStudy (anonymous):

really? i checked it in wolfram alpha

OpenStudy (jamesj):

Multiply out the expression I wrote down

OpenStudy (anonymous):

it becomes e^2t + 2e^(-2t) + e^-2t

OpenStudy (anonymous):

OH WAIT!

OpenStudy (jamesj):

No

OpenStudy (jamesj):

\[(e^t + e^{-t})^2 = e^{2t} + 2.e^t.e^{-t} + e^{-2t} = e^{2t} + 2 + e^{-2t}\]

OpenStudy (jamesj):

yes?

OpenStudy (anonymous):

oh jeez e^0 = 1

OpenStudy (anonymous):

thanks man!